OpenStudy (anonymous):

help i have 4 questions

OpenStudy (anonymous):

What is the approximate distance between the points (4, 5, 4) and (3, 7, 6)? 8 9 17<<<<<<<i think this is the answer. 29

OpenStudy (amistre64):

subtract, square, add, and sqrt is the process

OpenStudy (amistre64):

(4, 5, 4) -(3, 7, 6) ---------- 1 2 2 1+4+4 = 9 sqrt(9) = 3 hmm, thats not an option tho

OpenStudy (anonymous):

9 is

OpenStudy (amistre64):

make sure youre information is correctly posted

OpenStudy (anonymous):

yesp these are the choices 8 9 17 29

OpenStudy (anonymous):

wait oops its -4,5, 4 and 3,-7,-6

OpenStudy (anonymous):

oops im so sorry :(

OpenStudy (amistre64):

:) same process, different values

OpenStudy (amistre64):

-4,5, 4 -3,7,6 ---------- 7 12 10 49+144+100= 293 sqrt(293) is greater then 10, 11, 12, 13, 15, 17 is looking plausible yes

OpenStudy (anonymous):

so i was right 17?

OpenStudy (amistre64):

29^2 = 4xx 15^2 = 225 sqrt(293) = 17.11.... so yes

OpenStudy (anonymous):

cool okay here is the next one

OpenStudy (anonymous):

When the point (-1, 3) is dilated with the center of dilation at the origin, the image of the point is (-4, 12). What is the scale factor of this dilation? 4 0.25 - 4 -0.25

OpenStudy (amistre64):

-1 * s = -4 3 * s = 12 solve for s

OpenStudy (anonymous):

i get that part but i get confused when solving for s my original answer was 0.25 but i think im wrong b/c im unsure how to do it :(

OpenStudy (amistre64):

3, times what, equals 12? you have to know your basic times tables at least :)

OpenStudy (anonymous):

4 lol

OpenStudy (amistre64):

then the scale factor has to be 4 :)

OpenStudy (anonymous):

lol wow i was way way off lol thank you for your help on that one :P

OpenStudy (amistre64):

lol, thats ok, my first guess would have been 132

OpenStudy (anonymous):

lol ^.^ well the next two have pic that go with them .

OpenStudy (amistre64):

we gots a draw button underneath, or if you can do a screenshot and attach as a jpg thats fine

OpenStudy (anonymous):

Find the measure of APB if the measure of arc AB = 130°. 100° 70° 65° 50°

OpenStudy (amistre64):

theres a thrm that goes; the angle on the outside of what we are looking at is the difference between the outer arc and the inner arc

OpenStudy (anonymous):

lol thats a long term to remember haha :P

OpenStudy (amistre64):

they give the inner are as 130 the outter arc is 360 - 130 (360 - 130) - 130 should do it

OpenStudy (anonymous):

100 ^.^

OpenStudy (amistre64):

100 seems to work for me

OpenStudy (amistre64):

we can dbl chk another way tho

OpenStudy (amistre64):

the central angle is equal to the inner arc: 130 the tangents are always 90; so another 180 

OpenStudy (amistre64):

i guess i forgot the divide by 2 part of the thrm :/

OpenStudy (amistre64):

(outer-inner) ----------- 2

OpenStudy (anonymous):

Which pair of triangles is similar?

OpenStudy (anonymous):

okay sorry about that

OpenStudy (amistre64):

similar triangle have a common ratio of parts

OpenStudy (anonymous):

i thought it was number 1

OpenStudy (amistre64):

spose your lengths are: a,b,c and A,B,C a b c -- = -- = -- A B C

OpenStudy (anonymous):

ok

OpenStudy (amistre64):

number 1 is good

OpenStudy (anonymous):

but so is number 4 lol

OpenStudy (amistre64):

3 4 5 : 1 9 12 15 : 3

OpenStudy (amistre64):

number 4 is not

OpenStudy (amistre64):

3 4 5 -- = --- not= ---- 9 12 16

OpenStudy (anonymous):

really? i guess thats why i was confused i was stuck between the two

OpenStudy (amistre64):

helps if i use the right ratios .... #4 is squares

OpenStudy (amistre64):

3^2 = 9 4^2 = 16 5^2 = 25 these are not common rations\[\frac{5}{25}=\frac15\] \[\frac{4}{16}=\frac14\] \[\frac{3}{9}=\frac13\]

OpenStudy (anonymous):

oh i see and we found out that number 1 was 1/3 so the answer is number 1 ???

OpenStudy (amistre64):

correct

OpenStudy (anonymous):

thank you so much ^.^

OpenStudy (amistre64):

youre welcome, and good luck :)